OpenStudy (anonymous):

Eleven coins are tossed. What is the probability of getting at least one tail?

OpenStudy (anonymous):

one minus the probability that you get all heads

OpenStudy (anonymous):

how do i get the probability of getting heads?

OpenStudy (anonymous):

\(\frac{1}{2}\) for a head on one try \[\left(\frac{1}{2}\right)^{11}\] for heads eleven times in a row

OpenStudy (anonymous):

oh i see. thank you!